Circular No. 7937

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002eb IN CGCG 473-011 B. Beutler and W. D. Li, University of California at Berkeley, report

the LOTOSS (cf. IAUC 7906) discovery of an apparent supernova on unfiltered KAIT CCD images taken on July 22.4 (mag ∼ 18.5) and 23.4 UT (mag ∼ 18.2). SN 2002eb is located at α = 22^{h}19^{m}05^{s}.24, δ = +24^{o}35′39′′.8 (equinox 2000.0), which is 14′′.9 west and 13′′.7 south of the nucleus of CGCG 473-011. A KAIT image of the same field on July 15.4 UT showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2002ec IN NGC 5910 Beutler and Li also report the LOTOSS discovery of an apparent su-

pernova on unfiltered KAIT CCD images taken on July 23.2 (mag ∼ 17.7) and 24.2 UT (mag ∼ 17.5). SN 2002ec is located at α = 15^{h}19^{m}24^{s}.89, δ = +20^{o}53′14′′.0 (equinox 2000.0), which is 8′′.5 east and 12′′.9 south of the nucleus of the southwest component of the triple galaxy system NGC 5910. A KAIT image of the same field on Jul 15.2 UT showed nothing at this position (limiting mag ∼ 19.0).

NOVAE IN M31 Beutler and Li also report the LOTOSS discovery of two apparent

novae on unfiltered KAIT CCD images of M31. The first nova is located at α = 0^{h}43^{m}01^{s}.07, δ = +41^{o}16′20′′.1 (equinox 2000.0), which is 188′′.8 east and 11′′.6 north of the nucleus of M31. Unfiltered KAIT magnitudes for the first nova: Jan. 12.2 UT, [19.0; July 19.5, 18.3; 20.5, 18.0; 22.5, 17.9; 23.5, 18.1. The second nova is located at α = 0^{h}43^{m}06^{s}.51, δ = +41^{o}16′34′′.3 (equinox 2000.0), which is 250′′.1 east and 25′′.8 north of the nucleus of M31. Unfiltered KAIT magnitudes for the second nova: July 20.5 UT, [19.5; 22.5, 16.0; 23.5, 16.6.

COMET 57P/DU TOIT-NEUJMIN-DELPORTE Total magnitude estimates for the primary component (visual unless

otherwise noted): June 5.71 UT, 15.6: (A. Nakamura, Kuma, Ehime, Japan, 0.60-m reflector + CCD); 9.68, 15.4 (Nakamura); 18.70, 15.7 (K. Kadota, Ageo, Saitama, Japan, 0.18-m reflector + CCD); July 5.27, 13.4 (A. Hale, Cloudcroft, NM, 0.41-m reflector); 10.24, 13.3 (Hale); 10.95, 14.1 (K. Sárneczky, Ágasvár, Hungary, 0.38-m reflector); 11.59, 14.5 (Y. Ezaki, Toyonaka, Osaka, Japan, 0.30-m reflector + CCD).

2002 July 24 \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT Carl W. Hergenrother